# Generative modelling with jump-diffusions

### v1

@@ Abstract
Score-based diffusion models generate samples from an unknown target distribution using a time-reversed diffusion process. While such models represent state-of-the-art approaches in industrial applications such as artificial image generation, it has recently been noted that their performance can be further improved by considering injection noise with heavy tailed characteristics. Here, I present a generalization of generative diffusion processes to a wide class of non-Gaussian noise processes. I consider forward processes driven by standard Gaussian noise with super-imposed Poisson jumps representing a finite activity Lévy process. The generative process is shown to be governed by a generalized score function that depends on the jump amplitude distribution. Both probability flow ODE and SDE formulations are derived using basic technical effort, and are implemented for jump amplitudes drawn from a multivariate Laplace distribution. Remarkably, for the problem of capturing a heavy-tailed target distribution, the jump-diffusion Laplace model outperforms models driven by $\alpha$-stable noise despite not containing any heavy-tailed characteristics. The framework can be readily applied to other jump statistics that could further improve on the performance of standard diffusion models.
@@ 1 Introduction
Generative diffusion models transform a target distribution into a simple distribution by successively adding noise and then reversing this process. By sampling from the simple distribution, the reversed (or denoising) process is thus able to generate a new target sample representing, e.g., a new image of a tree. Starting from this seminal concept [15], two main implementations have been developed, where, roughly speaking, the denoising process is either formulated in discrete (time) steps and represented by a Markov chain, for example implemented in denoising score-matching (DSM) [19] and denoising diffusion probabilistic models (DDPMs) [15, 6], or implemented in the form of a stochastic differential equation (SDE) in continuous time [18]. In both cases, the underlying noise process is typically taken to have Gaussian characteristics, which facilitates both theoretical treatments and computational implementations.
@@
However, from a performance perspective, it is not a priori obvious that Gaussian noise represents the optimal choice in specific applications. Indeed, the question of how the performance of generative diffusion models can be improved by changing the characteristics of the underlying noise has recently attracted interest in the research community. In [11] it has been shown that using noise from a Gamma distribution in DDPM improves results for image and speech generation. In the SDE approach of [12], Gaussian noise has been replaced by noise from fractional Brownian motion, which achieves greater pixel-wise diversity and enhanced image quality. The effect of $\alpha$-stable noise has been investigated both in the SDE formulation [20] and DDPM [14], and has been shown to allow for faster and more diverse sampling. In [13] DDPM has been implemented with noise from a heavy-tailed Student-t distribution and shown to outperform standard diffusion models in heavy-tail estimation on high-resolution weather datasets. In [7] a Markovian jump model based on generator matching has been proposed, which achieves state-of-the-art performance in multi-modal protein generation.
In this work, I present a non-Gaussian generalization of the conventional SDE framework of diffusion models [18] that includes discontinuous jumps in the forward noising process. A similar idea has previously been explored in [20], but this approach is technically complicated and only tailored for a specific type of heavy-tailed noise. On the other hand, focusing on the mathematically much simpler class of jump-diffusion processes with finite activity, my framework represents a straightforward generalization of the Gaussian case and covers a wide-range of jump amplitudes.
The contributions of the present work are summarized as follows:
@@
  • A score-based generative model for jump-diffusions as underlying noise process is presented, generalizing the work of [18]
@@
  • Representations of the generative process in terms of a probability flow ODE and an SDE with jump-diffusion noise are derived. These are driven by a generalized score function $\neq\nabla\log\,p(\mathbf{x},t)$.
  • The detailed implementation for jump amplitudes drawn from a multivariate Laplace distribution is discussed. The model is shown to outperform methods based on heavy-tailed noise [14, 20] in capturing the tails of a heavy-tailed target distribution.
@@ 1.1 Relationship with other works
The present work is closely related to the work of [20], who first considered the case of jumps in the forward process in the Lévy-Ito model (LIM). While LIM only focuses on the special case of infinite activity $\alpha$-stable noise with $21<\alpha<$ 2, I consider finite activity Lévy noise given as a superposition of Gaussian noise with Poisson jumps drawn from an arbitrary normalizable distribution. The mathematical discussion is then much simplified and provides a pedagogical access to non-Gaussian generalizations of the conventional SDE framework. More details on the comparison with LIM are provided in Sec. C.2.
@@
In [7] a specific Markovian jump model has been proposed based on matching with a Gaussian probability path, i.e., the jump model generates the same conditional probability distributions as a Gaussian model, but has different sample paths. On the other hand, the finite activity Lévy model considered here is always characterized by non-Gaussian propagators.
@@ 2 Generative score-based diffusion models
for an initial data distribution $\mathbf{Y}(0)\sim p_{\rm data}(\mathbf{x})$ with sample space $\mathbf{x}\in\mathbb{R}^{d}$. The noise is typically taken as Gaussian white noise $\boldsymbol{\xi}(t)=\boldsymbol{\xi}^{\rm G}(t)$ with covariance $<\xi^{\rm G}_{i}(t)\xi^{\rm G}_{j}(t^{\prime})>=D\delta_{ij}\delta(t-t^{\prime})$. The forward process is distributed as $\mathbf{Y}(t)\sim p(\mathbf{x},t)$, with the probability density function (PDF) or propagator $p(\mathbf{x},t)=\int\mathrm{d}\mathbf{x}^{\prime}p(\mathbf{x},t|\mathbf{x}^{\prime})p_{\rm data}(\mathbf{x})$, where $p(\mathbf{x},t|\mathbf{x}^{\prime})$ denotes the conditional PDF of the forward process for the initial condition $\mathbf{Y}(0)=\mathbf{x}^{\prime}$.
@@
where the central ingredient is the score function $\nabla\log\,p(\mathbf{x},t)$, which directs the $\mathbf{X}$-process towards $p_{\rm data}$, and contains the PDF $p$ of the forward process. Describing the backwards evolution in terms of a PDF $\mathbf{X}(t)\sim\tilde{p}(\mathbf{x},t)$ and starting from $\tilde{p}(\mathbf{x},0)=p(\mathbf{x},T)$, we have $\tilde{p}(\mathbf{x},t)=p(\mathbf{x},T-t)$, i.e., the score function ensures that $\mathbf{X}$ at any time $t$ is equal in distribution to $\mathbf{Y}$ at time $T-t$. The challenge of sampling the starting point $\mathbf{X}(0)\sim p(\mathbf{x},T)$ is simplified by choosing $\mathbf{f},g$ such that $p(\mathbf{x},t|\mathbf{x}^{\prime})$ converges for $t\to\infty$ to a well-defined stationary distribution $p_{\rm st}(\mathbf{x})$, for example a simple Gaussian. For sufficiently large $T$, we have then $p(\mathbf{x},T)\approx p_{\rm st}(\mathbf{x})$ from which $\mathbf{X}(0)$ is sampled.
@@
$J_{\rm G}(\boldsymbol{\theta})$ $=$ $\int_{0}^{T}\mathrm{d}t\,w(t)\mathbb{E}_{\mathbf{Y}(0)}\mathbb{E}_{\mathbf{Y}(t)|\mathbf{Y}(0)}\left[\|\mathbf{s}_{\boldsymbol{\theta}}(\mathbf{Y},t)-\nabla\log\,p(\mathbf{Y},t|\mathbf{Y}_{0})\|^{2}\right]$ (4)
with weight function $w(t)$. Eq. (4) is called the denoising score matching loss function and highlights that the estimation of $\mathbf{s}_{\boldsymbol{\theta}}$ does not require knowledge of the exact $p_{\rm data}$, but can be performed from finite samples of $p_{\rm data}$ using Monte-Carlo integration.
@@ 3 From diffusions to jump-diffusions
The aim of this work is to consider the case when the noise $\boldsymbol{\xi}$ in Eq. (1) is not only given as Gaussian white noise, but also contains non-Gaussian contributions in the form of additional discontinuous jumps that occur with a fixed rate and whose amplitudes are drawn independently from a given distribution. The general class of such processes with stationary and independent increments is known as Lévy processes and have a well-developed mathematical theory [4]. In the following, I focus on the class of finite activity Lévy processes, for which the jump amplitude distribution is normalizable, excluding the case of power-law or gamma distributed jump amplitudes. The mathematical theory is considerably simpler in this case and leads to a straightforward generalization of the conventional Gaussian case discussed in Sec. 2. The special case of pure power-law jumps corresponding to $\alpha$-stable noise with $21<\alpha<$ 2 has been discussed in detail in the LIM framework [20] using rigorous methods, see also Sec. C.2.
@@
$G_{\Delta\boldsymbol{\xi}}(\mathbf{k})=\left<e^{i\mathbf{k}{\cdot}\Delta\boldsymbol{\xi}}\right>=e^{-\psi(k)\Delta t},\qquad\qquad\psi(k)=\frac{D}{2}k^{2}-\lambda\phi(k)$ (7)
@@
$\phi(k)=\int\mathrm{d}\mathbf{z}\left(e^{i\mathbf{k}{\cdot}\mathbf{z}}-1\right)\rho(z).$ (8)
contains the characteristic function of the jumps. A crucial constraint is that I assume isotropic jump distributions such that $\rho$ is a function of $z=\|\mathbf{z}\|$ only and, as a consequence, also $\phi$ is a function of $k=\|\mathbf{k}\|$ only.
I then obtain the following jump-diffusion generalizations of Eqs. (2,3,4), when the noise of Eq. (1) is specified by Eqs. (7,8):
@@
    Generalized score function: Instead of $\nabla\log\,p(\mathbf{x},t)$, the score function that governs the generative process is now given by $\frac{\nabla\mathcal{V}(\mathbf{x},t)}{p(\mathbf{x},t)}$, where $p(\mathbf{x},t)$ is the PDF of the forward process $\mathbf{Y}(t)$ with Fourier transform (FT) $\hat{p}(\mathbf{k},t)$ and $\mathcal{V}$ is expressed in terms of the inverse FT
@@
    Generative process: The jump-diffusion equivalent of the probability flow equation Eq. (3) is given by
@@
    $\dot{\mathbf{X}}(t)=-\mathbf{f}(\mathbf{X},T-t)+2\frac{\nabla\mathcal{V}(\mathbf{X},T-t)}{p(\mathbf{X},T-t)}+g(T-t)\boldsymbol{\xi}(t)$ (11)
    For the derivation of Eqs. (9,10,11) see Appendix A
@@
    $J_{\rm JD}(\boldsymbol{\theta})$ $=$ $\int_{0}^{T}\mathrm{d}t\,w(t)\mathbb{E}_{\mathbf{Y}(0)}\mathbb{E}_{\mathbf{Y}(t)|\mathbf{Y}(0)}\left[\left\|\mathbf{s}_{\boldsymbol{\theta}}(\mathbf{Y},t)-\frac{\nabla\mathcal{V}(\mathbf{Y},t|\mathbf{Y}_{0})}{p(\mathbf{Y},t|\mathbf{Y}_{0})}\right\|^{2}\right]$ (12)
    where
@@
    and $p(\mathbf{x},t|\mathbf{x}^{\prime})$ denotes the propagator of $\mathbf{Y}(t)$, Eq. (1), with FT $\hat{p}(\mathbf{k},t|\mathbf{x}^{\prime})$. Eq. (12) can be derived from the explicit score matching loss function following standard steps, see Appendix B
@@
    where $\mathbf{s}_{\boldsymbol{\theta}}$ is estimated by Eq. (12) and $\mathbf{Z}_{i}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. The jump increment $\Delta\mathbf{J}_{i}$ is generated by drawing $M\sim\text{Poisson}(\lambda\Delta t)$ and setting $\Delta\mathbf{J}_{i}=\sum_{j=1}^{M}\mathbf{A}_{j}$, where $\mathbf{A}_{j}\sim\rho$ i.i.d. One obtains likewise a jump-diffusion variant of DDPM by setting $\mathbf{f}(\mathbf{x},t)=-\frac{1}{2}\beta(t)\mathbf{x}$, $g(t)=\sqrt{\beta(t)}$.
@@
3.1 The jump-diffusion Laplace model
In the Gaussian case, it is desirable to choose $\mathbf{f},g$ such that: (i) a closed analytical form of the propagator $p(\mathbf{x},t|\mathbf{x}^{\prime})$ is available, which facilitates the score function approximation via Eq. (4); and (ii) to be able to generate samples from $p_{\rm st}(\mathbf{x})$ efficiently. A common choice is thus an Ornstein-Uhlenbeck process as forward process, where $p(\mathbf{x},t|\mathbf{x}^{\prime})$ is Gaussian. In the jump-diffusion case, we need to consider the loss function Eq. (12) and thus also need to evaluate the term $\nabla\mathcal{V}(\mathbf{x},t|\mathbf{x}^{\prime})$. As we will see, the expressions are more involved as in the Gaussian case, but remain computationally tractable due to the restriction to isotropic jump amplitudes.
In the detailed implementation, I will consider $\mathbf{Y}(t)$ as a (generalized) Ornstein-Uhlenbeck process with $\mathbf{f}(\mathbf{x},t)=\frac{1}{2}\mathbf{x}$ and $g(t)=1$ such that Eq. (1) becomes
$\dot{\mathbf{Y}}(t)=-\frac{1}{2}\mathbf{Y}(t)+\boldsymbol{\xi}^{\rm G}(t)+\sum_{j=1}^{N_{T}}\mathbf{A}_{j}\delta(t-\tau_{j})$ (15)
In addition, I consider jumps with amplitudes $\mathbf{A}_{j}$ drawn from an isotropic multivariate Laplace distribution: $\mathbf{A}_{j}\sim\text{GL}(\sigma^{2},1)$, see Appendix D. This distribution not only represents a multivariate extension of the exponential distribution and is thus a standard case to consider beyond the Gaussian, but it also allows for straightforward sampling and has a simple characteristic function $G_{\rm L}(k)=(1+\frac{\sigma^{2}}{2}k^{2})^{-1}$, which simplifies the solution of the Ornstein-Uhlenbeck process Eq. (15). I refer to the resulting model as jump-diffusion Laplace (JDL) model. The propagator is then obtained by solving the Fokker-Planck equation associated with Eq. (15) leading to the solution (see Appendix C)
$p(\mathbf{x},t|\mathbf{x}^{\prime})$ $=$ $G_{1}(|\mathbf{x}-\mathbf{x}^{\prime}e^{-t/2}|,t)$ (16)
where $G_{1}(x,t)$ is given in terms of the inverse FT $G_{1}(x,t)=\mathcal{F}^{-1}\left\{\hat{G}_{1}(k,t)\right\}$ with
$\hat{G}_{1}(k,t)$ $=$ $\left(\frac{1+\frac{\sigma^{2}}{2}k^{2}e^{-t}}{1+\frac{\sigma^{2}}{2}k^{2}}\right)^{\lambda}e^{-\frac{D}{2}k^{2}\left(1-e^{-t}\right)}.$ (17)
Since $\hat{G}_{1}(k,t)$ is isotropic in $k$, evaluating $G_{1}(x,t)$ only requires a one-dimensional integral for arbitrary $d$. Likewise, the term $\nabla\mathcal{V}(\mathbf{x},t|\mathbf{x}^{\prime})$ in Eq. (12) can be evaluated by a one-dimensional integral due to isotropy
$\nabla\mathcal{V}(\mathbf{x},t|\mathbf{x}^{\prime})=\frac{\mathbf{x}-\mathbf{x}^{\prime}e^{-t/2}}{|\mathbf{x}-\mathbf{x}^{\prime}e^{-t/2}|}G_{2}(|\mathbf{x}-\mathbf{x}^{\prime}e^{-t/2}|,t)$ (18)

### v2

@@ Abstract
Score-based diffusion models generate samples from an unknown target distribution using a time-reversed diffusion process. While such models represent state-of-the-art approaches in industrial applications such as artificial image generation, it has recently been noted that their performance can be further improved by considering injection noise with heavy tailed characteristics. Here, I present a generalization of generative diffusion processes to a wide class of non-Gaussian noise processes. I consider forward processes driven by standard Gaussian noise with super-imposed Poisson jumps representing a finite activity Lévy process. The generative process is shown to be governed by a generalized score function that depends on the jump amplitude distribution and can be estimated by minimizing a simple MSE loss as in conventional Gaussian models. Both probability flow ODE and SDE formulations are derived using basic technical effort. A detailed implementation for a pure jump process with Laplace distributed amplitudes yields a generalized score function in closed analytical form and is shown to outperform the equivalent Gaussian model in specific parameter regimes.
@@ 1 Introduction
Generative diffusion models transform a target distribution into a simple distribution by successively adding noise and then reversing this process. By sampling from the simple distribution, the reversed (or denoising) process is thus able to generate a new target sample. Starting from this seminal concept [21], two main implementations of diffusion models (DMs) have been developed, where, roughly speaking, the denoising process is either formulated in discrete (time) steps and represented by a Markov chain, as implemented for example in denoising score-matching [26] and denoising diffusion probabilistic models (DDPMs) [21, 8], or implemented in the form of an ordinary differential equation (ODE) or stochastic differential equation (SDE) in continuous time [24]. DMs based on ODEs are closely related to optimal transport problems and have been generalized in frameworks known as flow matching [13, 14, 25] and stochastic interpolants [1].
@@
The overwhelming majority of research on DMs is based on underlying noise processes or distributions with Gaussian characteristics, which facilitates both theoretical treatments and computational implementations. However, from a performance perspective, it is not a priori obvious that Gaussian statistics represents the optimal choice in specific applications. Indeed, the question of how the performance of generative diffusion models can be improved by changing the characteristics of the underlying noise has recently attracted interest in the research community. In [15] it has been shown that using noise from a Gamma distribution in DDPM improves results for image and speech generation. In the SDE approach of [16], Gaussian noise has been replaced by noise from fractional Brownian motion, which achieves greater pixel-wise diversity and enhanced image quality. The effect of $\alpha$-stable noise has been investigated both in the SDE formulation [27] and DDPM [20], and has been shown to allow for faster and more diverse sampling. In [17] DDPM has been implemented with noise from a heavy-tailed Student-t distribution and shown to outperform standard diffusion models in heavy-tail estimation on high-resolution weather datasets. In [9] a Markovian jump model based on generator matching has been proposed, which achieves state-of-the-art performance in multi-modal protein generation.
In this work, I present a non-Gaussian generalization of the conventional Gaussian framework of diffusion models that includes discontinuous jumps in the noise process. The contributions of the present work are summarized as follows:
@@
  • A score-based generative framework for jump-diffusions as underlying noise process is presented, generalizing the work of [24].
  • The framework is based on a generalized score function $\neq\nabla\log\,p(\mathbf{x},t)$ that can be estimated by minimizing a simple denoising score matching loss function similar to the standard case [26].
@@
  • Representations of the generative process in terms of a probability flow ODE and an SDE with jump-diffusion noise are derived.
  • A detailed implementation for jump amplitudes drawn from a multivariate Laplace distribution is presented (referred to as JL model), where both conditional distributions and the generalized score function have closed analytical forms. The training and generative sampling is only slightly more computationally costly than for standard DMs and is shown to outperform an equivalent Gaussian model in specific parameter regimes.
@@ 1.1 Relationship with other works
DMs driven by non-Gaussian noise. The present work is closely related to the work of [27], who first considered the case of jumps in the forward process in the Lévy-Ito model (LIM). While LIM only focuses on the special case of infinite activity $\alpha$-stable noise with $1<\alpha<2$, I consider finite activity Lévy noise given as a superposition of Gaussian noise with Poisson jumps drawn from an arbitrary normalizable distribution. The mathematical discussion is then much simplified and provides a pedagogical access to non-Gaussian generalizations of the conventional framework. Moreover, the LIM SDE neglects an additional noise term with finite variation, which arises from the exact time reversal formalism [5, 27]. As such noise and score-function are not exactly balanced unlike in the JL model presented here. The computational cost of training and sampling the JL model is equivalent to LIM and thus provides an interesting non-Gaussian alternative to consider in specific applications. For the simple tests and neural network architecture considered here the JL model is shown to outperform LIM in a range of parameter regimes.
@@
DMs based on general Markov processes. Recently, frameworks for DMs have been presented that include jump processes on continuous state spaces as well [9, 18]. The “generator-matching" framework of [9] specifies the backward generative process in terms of the generator of a general Markov process and learns its coefficients using a prescribed probability path, thus generalizing the flow matching approach to general SDEs. Extending ideas formulated in [3], the denoising framework of [18] considers a general Markov process as forward process and estimates the corresponding backward process by minimizing a KL divergence as loss function. However, applying this approach to finite activity Lévy processes generally leads to backward processes with jump distributions that are inhomogeneous in space and time and thus non-trivial to sample from. Likewise in generator-matching, where the backwards generator depends on the prescribed probability path, but contains space- and/or time-inhomogeneous jump distributions already for very simple paths. On the other hand, the generalized score function derived here is shown to balance the noise in the backward process and can be estimated using a simple MSE loss. Instead of estimating the backward jump distribution as in [9, 18], the framework thus only requires the estimation of a drift term, which captures both ODE and SDE representations. As such it represents an elegant and simple generalization of the Gaussian case.
@@ 2 Generative score-based diffusion models
for an initial data distribution $\mathbf{Y}(0)\sim p_{\rm data}(\mathbf{x})$ with sample space $\mathbf{x}\in\mathbb{R}^{d}$. The noise is typically taken as Gaussian white noise $\boldsymbol{\xi}(t)=\boldsymbol{\xi}^{\rm G}(t)$ with covariance $\xi^{\rm G}_{i}(t)\xi^{\rm G}_{j}(t^{\prime})>=D\delta_{ij}\delta(t-t^{\prime})$. The forward process is distributed as $\mathbf{Y}(t)\sim p(\mathbf{x},t)$, with the associated probability density function (PDF) $p(\mathbf{x},t)=\int\mathrm{d}\mathbf{x}^{\prime}p(\mathbf{x},t|\mathbf{x}^{\prime})p_{\rm data}(\mathbf{x})$, where $p(\mathbf{x},t|\mathbf{x}^{\prime})$ denotes the conditional PDF of the forward process for the initial condition $\mathbf{Y}(0)=\mathbf{x}^{\prime}$.
@@
where the central ingredient is the score function $\nabla\log\,p(\mathbf{x},t)$, which directs the $\mathbf{X}$-process towards $p_{\rm data}$, and contains the PDF $p$ of the forward process. Describing the backwards evolution in terms of a PDF $\mathbf{X}(t)\sim\tilde{p}(\mathbf{x},t)$ and starting from $\tilde{p}(\mathbf{x},0)=p(\mathbf{x},T)$, we have $\tilde{p}(\mathbf{x},t)=p(\mathbf{x},T-t)$, i.e., the score function ensures that $\mathbf{X}$ at any time $t$ is equal in distribution to $\mathbf{Y}$ at time TtT-t. The challenge of sampling the starting point $\mathbf{X}(0)\sim p(\mathbf{x},T)$ is simplified by choosing $\mathbf{f},g$ such that $p(\mathbf{x},t|\mathbf{x}^{\prime})$ converges for $t\to\infty$ to a well-defined stationary distribution $p_{\rm st}(\mathbf{x})$, for example a simple Gaussian. For sufficiently large $T$, we have then $p(\mathbf{x},T)\approx p_{\rm st}(\mathbf{x})$ from which $\mathbf{X}(0)$ is sampled.
@@
$J_{\rm G}(\boldsymbol{\theta})$ $=$ $\int_{0}^{T}\mathrm{d}t\,w(t)\mathbb{E}_{\mathbf{Y}_{0}\sim p_{\rm data}}\mathbb{E}_{\mathbf{Y}(t)|\mathbf{Y}_{0}}\left[\|\mathbf{s}_{\boldsymbol{\theta}}(\mathbf{Y},t)-\nabla\log\,p(\mathbf{Y},t|\mathbf{Y}_{0}\|^{2}\right]$ (4)
with weight function $w(t)$. Eq. (4) is called the denoising score matching (DSM) loss function [26] and highlights that the estimation of $\mathbf{s}_{\boldsymbol{\theta}}$ does not require knowledge of the exact $p_{\rm data}$, but can be performed from finite samples of $p_{\rm data}$ using Monte-Carlo integration.
@@ 3 From diffusions to jump-diffusions
The aim of this work is to consider the case when the noise $\boldsymbol{\xi}$ in Eq. (1) is not given as purely Gaussian white noise, but contains non-Gaussian contributions in the form of discontinuous jumps that occur with a fixed rate and whose amplitudes are drawn independently from a given distribution. The general class of such processes with stationary and independent increments is known as Lévy processes and have a well-developed mathematical theory [6]. In the following, I focus on the class of finite activity Lévy processes, for which the jump amplitude distribution is normalizable, excluding the case of power-law or gamma distributed jump amplitudes. The mathematical theory is considerably simpler in this case and leads to a straightforward generalization of the conventional Gaussian case discussed in Sec. 2. The special case of pure power-law jumps corresponding to $\alpha$-stable noise with $1<\alpha\leq 2$ has been discussed in detail in the LIM framework [27] using rigorous methods.
@@
$G_{\Delta\boldsymbol{\xi}}(\mathbf{k})=\mathbb{E}\left[e^{i\mathbf{k}{\cdot}\Delta\boldsymbol{\xi}}\right]=e^{-\psi(k)\Delta t},\qquad\qquad\psi(k)=\frac{D}{2}k^{2}-\lambda\phi(k)$ (7)
@@
$\phi(k)=\int\mathrm{d}\mathbf{z}\left(e^{i\mathbf{k}{\cdot}\mathbf{z}}-1\right)\rho(\mathbf{z}).$ (8)
contains the characteristic function of the jumps. A crucial constraint is that I assume isotropic jump distributions such that $\rho$ is a function of $z=\|\mathbf{z}\|$ only and, as a consequence, $\phi$ can be expressed as a function of $k=\|\mathbf{k}\|$ only. Moreover, $\rho$ needs to be normalizable: $\int\mathrm{d}\mathbf{z}\,\rho(\mathbf{z})=1$.
The following jump-diffusion generalization of Eqs. (2,3,4) is obtained when the noise of Eq. (1) is specified by Eqs. (7,8):
@@
    Generalized score function: Instead of $\nabla\log\,p(\mathbf{x},t)$, the score function that governs the generative process is now given by
    ${\mathcal{S}(\mathbf{x},t)=\frac{\nabla\mathcal{V}(\mathbf{x},t)}{p(\mathbf{x},t)}}$ (9)
    where $p(\mathbf{x},t)$ is the PDF of the forward process $\mathbf{Y}(t)$ with Fourier transform (FT) $\hat{p}(\mathbf{k},t)$ and $\mathcal{V}$ is expressed in terms of the inverse FT
@@
    Generative process: The jump-diffusion equivalent of the probability flow ODE Eq. (3) is given by
    ${\dot{\mathbf{X}}(t)=-\mathbf{f}(\mathbf{X},T-t)+\mathcal{S}(\mathbf{X},T-t)}$ (11)
@@
    ${\dot{\mathbf{X}}(t)=-\mathbf{f}(\mathbf{X},T-t)+2\,\mathcal{S}(\mathbf{X},T-t)+g(T-t)\boldsymbol{\xi}(t)}$ (12)
    For the derivation of Eqs. (10,11,12) see Appendix A
@@
    $J_{\rm JD}(\boldsymbol{\theta})$ $=$ $\int_{0}^{T}\mathrm{d}t\,w(t)\mathbb{E}_{\mathbf{Y}_{0}\sim p_{\rm data}}\mathbb{E}_{\mathbf{Y}(t)|\mathbf{Y}_{0}}\left[\left\|\mathbf{s}_{\boldsymbol{\theta}}(\mathbf{Y},t)-{\mathcal{S}(\mathbf{Y},t|\mathbf{Y}_{0})}\right\|^{2}\right]$ (13)
    Here, I introduce the conditional generalized score function $\mathcal{S}(\mathbf{x},t|\mathbf{x}^{\prime})=\frac{\nabla\mathcal{V}(\mathbf{x},t|\mathbf{x}^{\prime})}{p(\mathbf{x},t|\mathbf{x}^{\prime})}$, where
@@
    and $p(\mathbf{x},t|\mathbf{x}^{\prime})$ denotes the conditional PDF of $\mathbf{Y}(t)$, Eq. (1), with FT $\hat{p}(\mathbf{k},t|\mathbf{x}^{\prime})$. Eq. (13) can be derived from the explicit score matching loss function following standard steps, see Appendix B
@@
    where $\mathbf{s}_{\boldsymbol{\theta}}$ is estimated by Eq. (13) and $\mathbf{Z}_{i}\sim\mathcal{N}(\mathbf{0},D\mathbf{I})$. The jump increment $\Delta\mathbf{J}_{i}$ is generated by drawing $M\sim\text{Poisson}(\lambda\Delta t)$ and setting $\Delta\mathbf{J}_{i}=\sum_{j=1}^{M}\mathbf{A}_{j}$, where $\mathbf{A}_{j}\sim\rho$ i.i.d. One obtains likewise a jump-diffusion variant of DDPM by setting $\mathbf{f}(\mathbf{x},t)=-\frac{1}{2}\beta(t)\mathbf{x}$, $g(t)=\sqrt{\beta(t)}$.
@@
3.1 The JL model: Laplace distributed jump amplitudes
For efficient generative modelling, it is desirable to have: (i) a closed analytical form of the conditional score function, which facilitates the loss function minimization; and (ii) to be able to generate samples from $p_{\rm st}(\mathbf{x})$ efficiently. For standard DMs, a common choice is thus an Ornstein-Uhlenbeck process as forward process, where $p(\mathbf{x},t|\mathbf{x}^{\prime})$ is a Gaussian and thus satisfies both requirements. In the jump-diffusion case, I consider likewise $\mathbf{Y}(t)$ as a (generalized) Ornstein-Uhlenbeck process, where for simplicity
$\mathbf{f}(\mathbf{x},t)=-\frac{1}{2}\,\mathbf{x}\qquad,\qquad\qquad g(t)=1$ (16)
The choice of amplitude distribution of course significantly affects the conditional PDF and generalized score function. The non-Gaussian model becomes particularly simple for the case of a pure jump process ($D=0$) with $\lambda=1$ and amplitudes drawn from an isotropic multivariate Laplace distribution: $\mathbf{A}_{j}\sim\text{L}_{d}(\sigma^{2})$. The distribution $\text{L}_{d}(\sigma^{2})$ not only represents a multivariate extension of the exponential distribution and is thus a standard case to consider beyond the Gaussian, but also allows for straightforward sampling. The resulting diffusion model is given in closed analytical form and is only slightly more computationally costly than the standard Gaussian case. I refer to this model as jump Laplace (JL) model.
For details of the derivation I refer to Appendix C. The conditional PDF of the JL model is given by
$p(\mathbf{x},t|\mathbf{x}^{\prime})$ $=$ $\rho_{\rm L}(\mathbf{x}-\mathbf{x}^{\prime}e^{-t/2},\sigma)\left(1-e^{-t}\right)+\delta(\mathbf{x}-\mathbf{x}^{\prime}e^{-t/2})e^{-t}$ (17)